Find two consecutive odd numbers whose product is 195.
Answer by oberobic(2304)

Two consecutive odd integers can be defined as: x and y.
A better way to define y is to say it = x+2, which means you have only one unknown.
We're told x*y = 195, so we can substitute y=x+2.
.
x(x+2) = 195
x^2 + 2x = 195
x^2 + 2x - 195 = 0
Factoring...
(x + 15)(x - 13) = 0
.
So, x can = -15 or 13.
.
Since the question does not say "positive consecutive odd numbers", we need to check both to ensure they're both solutions.
.
x = -15
y = x+2 = -13
(-13)*(-15) = 195
.
x = 13
y = x+2 = 15
13*15 = 195
.
Answer:
Two consecutive odd numbers whose product is 195 is either 13 & 15 or -15 & -13.